OpenStudy (anonymous):

Could someone help me with this? will fan. A can of spaghetti has a diameter of 12 cm and a height of 16 cm. The can contains spherical meatballs, each with a radius of 3 cm, that take up 25% of the volume. How many meatballs are in the can? Use 3.14 to approximate pi and express your answer as a whole number.

OpenStudy (anonymous):

vol.of tin =pi *r^2*h vol.of sphere =4/3*pi*r^3 take 25% of vol of tin =vol of spheres divide by vol of each sphere gives number of spheres

OpenStudy (anonymous):

so once i find the volume for both, i divide the volume of the can by the volume of the spheres?

OpenStudy (anonymous):

once you take 25% of the tin vol as the question says which of course is a 1/4 that = spheres then divide by vol. of sphere to get how many

OpenStudy (anonymous):

okay, so once I do the 25%, would the 25% be 452.16? Then if it is, 452.16/113.04 = 4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, thank you for your help!! :)